[Brahms-l] Shift report 20050529 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun May 29 2005 - 07:57:36 EDT
Shifter leader was Michael Murray
Shifters were Theresa Wilkerson


1. Beam normal throughout shift.

2. Ended run #15036. Runs taken on shift were #15037-15052.

3. Problems with BBL and BBR voltages. Also CCL voltage on #10 and 17. Several attempts made to resolve; ended shift with BB ok, but they did not stay that way earlier in shift after fixes.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun May 29 07:57:50 2005

This archive was generated by hypermail 2.1.8 : Sun May 29 2005 - 07:58:03 EDT